let T:P2-->P2 defined by T(a+bx+cx^2)=(2b+3cx+(a-b)x^2) check T is linear trnsfrmtn.find matrix wrt bases B={x2,x2+x,x2+x+1} ,B2={1,x,x2}.find kernel of T
1
Expert's answer
2020-04-15T17:04:34-0400

"T(a+bx+cx^2)=2b+3cx+(a-b)x^2"

It should satisfy

"(i)T(u+v)=T(u)+T(v)"

"(ii)T(mu)=mT(u)"

"u=f+ex+dx^2;v=i+hx+gx^2"


"(i)T(u+v)=T((i+f)+(e+h)x+(d+g)x^2)"

"=2(e+h)+3(d+g)x+(i+f-c-h)x^2"

"=[2e+3dx+(f-e)x^2]+[2h+3gx+(i-h)x^2]"

"=T(u)+T(v)"


"(ii)T(mu)=T(mf+mex+mdx^2)"

"=2me+3mdx+(mf-me)x^2"

"=m(2e+3dx+(f-e)x^2)"

"=mT(u)"


So,T is a linear Transformation.


"B_1=({x^2,x^2+x,x^2+x+1})"

"T(x^2)=3x=3(x^2+x)-3(x^2)+0.(x^2+x+1)"

"T(x^2+x)=2+3x-x^2=(-4).x^2+1.(x^2+x)+2.(x^2+x+1)"

"T(x^2+x+1)=2+3x=(-3).x^2+1.(x^2+x)+2.(x^2+x+1)"

The coefficient of T(x2) is "\begin{pmatrix}
 -3 \\
 3\\
0
\end{pmatrix}"


The coefficient of T(x2+x) is "\begin{pmatrix}
 -4 \\
 1\\
2
\end{pmatrix}"


The coefficient of T(x2+x+1) is "\begin{pmatrix}
 -3 \\
 1\\
2
\end{pmatrix}"


So the matrix is "\begin{bmatrix}
 -3& -4 &-3\\
 3 & 1&1\\
0&2&2
\end{bmatrix}"

"B_2=(1,x,x^2)"

"T(1)=x^2"

"T(x)=2-x^2"

"T(x^2)=3x"

The coefficient of T(1) is "\begin{pmatrix}
 0 \\
 0\\
1
\end{pmatrix}"


The coefficient of T(x) is "\begin{pmatrix}
 2 \\
 0\\
-1
\end{pmatrix}"


The coefficient of T(x2) is "\begin{pmatrix}
 0 \\
 3\\
0
\end{pmatrix}"


So,the matrix is "\begin{bmatrix}
 0 & 2 &0\\
 0 & 0&3\\
1&-1&0
\end{bmatrix}"


Ker T={v: "T(v)=0" }

"2b+3cx+(a-b)x^2=0"

"2b=0;3c=0;(a-b)=0"

"b=0,c=0,a=b=0"

"a=b=c=0"

So,v=0

Ker T={0}